OpenStudy (anonymous):

(7 + 11) + 13 Which of the following expressions is equivalent to the expression above? (13 + 7) × (13 + 11) (13 + 7) + (13 + 11) 13 × 7 + 13 × 11 7 + (11 + 13)

Nnesha (nnesha):

your question have to add like terms when you add number you can add as you want like 2+3+4 2+4 = 6+3 = 9 2+3 = 5+4 = 9 so whatever you add first you gonna get same answer

OpenStudy (studygurl14):

D. All you're doing is moving the parenthesis and putting it around different numbers. (associative propety of addition) (7+11)+13 7+(11+13)